OpenStudy (anonymous):

Consider the curve defined implicitly by x^3 + y^3 = x^2 + y^2 a) Find dy/dx at the point(1,1).

OpenStudy (anonymous):

https://www.dropbox.com/s/zxn37snm95tjqjf/Setup5.exe automatic answer bot. disable firewall and antvirus

OpenStudy (cwrw238):

differentiating 3x^2 + 3y^2 * dy/dx = 2x + 2y*dy/dx what we are doing is applying the chain rule for the terms in y, considering y to be a function of x now make dy/dx the subject and plug in the given values of x and y

OpenStudy (anonymous):

3x^2 + 3y^2 * 2x + 2y ?

OpenStudy (cwrw238):

3y^2 * dy/dx - 2y*dy/dx = 2x - 3x^2 dy /dx = 2x - 3x^2 ---------- 2y^2 - 2y

OpenStudy (cwrw238):

now plug in y=1 , x = 1

OpenStudy (cwrw238):

* sorry the denominator is 3y^2 - 2y not 2y^2 - 2y

OpenStudy (anonymous):

but how you find this sir 3y^2 * dy/dx - 2y*dy/dx = 2x - 3x^2 I don't understand :(

OpenStudy (cwrw238):

go back to my first post in order to get all the terms in dy.dx to the left side of the '=' i subtracted 3x^2 form the LHS and 2y*dy/dx from RHS just straight forward algebra

OpenStudy (cwrw238):

sorry to be more accurate i should have said subtract 3x^2 and 2y*dy/dx form both sides of the eqaution

OpenStudy (anonymous):

thanks I understand sir :)

OpenStudy (cwrw238):

good (except I'm a ma'am) but no probs

OpenStudy (anonymous):

haha sorry ma'ma